OpenStudy (anonymous):

- 9 (r + 5) = -4r – 10

OpenStudy (anonymous):

-9r-45=-4r-10 -5r=45-10 -5r=35 r=-7

OpenStudy (anonymous):

-5(3x + 5) = -10x + 20 is problem also @koymoi

OpenStudy (anonymous):

-5(3x + 5) = -10x + 20 -15x - 25 = -10x + 20 -15x -25+25 = -10x + 20 + 25 -15x = -10x + 45 -15x + 10x = -10x + 10x + 45 -5x = 45 -5x/-5 = 45/-5 x = -9 I think

OpenStudy (mathlegend):

@koymoi and @LoveYou*69 I would like to see you not just give the answer basically. Let the OP do the actual work. Let the OP think and let you see what step is next. Ask the OP questions and explain what you are doing please. :D

OpenStudy (skullpatrol):

I agree, doing all the algebraic steps for a person is not going to help them on a test.

OpenStudy (anonymous):

@MathLegend okay, sorry, im new to this site. heh

OpenStudy (skullpatrol):

OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.